OpenStudy (anonymous):

Thor invests $750 in an account, with interest compounded continuously. If his investment doubles in value after 9 years, how much money will he have after 12 years? [Use the formla A = Pert and round answer to the nearest cent.]

OpenStudy (anonymous):

u have P = 750 t=9 A = 2*750 = 1500 then use A=Pe^(rt) to solve for r and then use A=Pe^(rt) but with t=12